# Transitioning from perturbative to non-perturbative splittings within QCD jets

### v2

@@ Abstract
The study of Quantum Chromodynamics (QCD) at ultra-relativistic energies can be performed in a controlled environment through lepton-hadron deep inelastic scatterings. In such collisions, the high-energy partonic emissions that follow from the ejected hard partons are accurately described by perturbative QCD. However, the lower energy scales at which quarks and gluons experience colour confinement, i.e. hadronization mechanism, fall outside the validity regions for perturbative calculations, requiring phenomenological models tuned to data to describe it. As such, hadronization physics cannot be currently derived from first principles alone. Monte Carlo event generators are useful tools to describe these processes as they simulate both the perturbative and the non-perturbative interactions, with model-dependent energy scales that control parton dynamics. This work aims to use jets - experimental reconstructions of final-state particles likely to have a common partonic origin - to inspect this transition further. Although originally proposed to circumvent hadronization effects, we show that jets can be utilised as probes of non-perturbative phenomena via their substructure. The charge correlation ratio was recently shown to be sensitive to hadronization effects. Our work further improves this sensitivity to non-perturbative scales by introducing a new selection based on the relative placement of the resolved SoftDrop splitting within the clustering tree, defined as the unclustering that resolves the jet’s leading charged particles.
@@ I I - Introduction
Recent effort in the high-energy physics community has been in the area of developing novel experimental algorithms that translate a jet clustering tree to a theoretically motivated description of a parton shower Dasgupta and Salam (2001); Dasgupta et al. (2013); Larkoski et al. (2014); Mehtar-Tani et al. (2020). Measurements from the Relativistic Heavy-Ion Collider (RHIC) and from the Large Hadron Collider (LHC) have shown that jet substructure observables, when calculated via grooming techniques such as SoftDrop (SD) Sirunyan et al. (2018a, b); Aad et al. (2020a); Acharya et al. (2020); Aad et al. (2020b); Adam et al. (2020); Abdallah et al. (2021); Acharya et al. (2022); ALICE (2022a, b), allow one to fit pQCD calculations at the first hard splitting without the need for large hadronization corrections. Namely, the groomed momentum fraction ($z_{g}$), i.e. the fractional transverse momentum carried by the softest emitted parton, fits very well to the pQCD splitting functions when evaluated at the first SD splitting. For subsequent emissions along the higher energy (hard) branch of a jet shower Dreyer et al. (2018), the shape of the splitting function gradually changes from a steeply falling pQCD-like $1/z$ distribution to a flatter shape due to phase space restrictions. However, so as the impact of npQCD effects grows Robotkova (2022); Kunnawalkam Elayavalli (2022). Previous works Apolinário et al. (2021) demonstrated that SD groomed clustering trees significantly improve the correlation between the formation time of the parton shower emissions and the formation time of the first jet SD unclustering step, regardless of choice of jet clustering algorithm Apolinário et al. (2021). However, subsequent unclustering steps also observe their parton shower correlation worsen, although with a large dependence on the chosen algorithm.
@@
The current state of SoftDrop substructure studies therefore requires the assessment of non-perturbative contributions to the jet evolution history, as they become more and more meaningful to the late-time jet dynamics. This calls for the exploration of experimentally robust jet substructure observables sensitive to hadronization physics. In Chien et al. (2022) the charge correlation ratio was proposed as a possible observable that is able to discriminate between two different classes of hadronization models: the Lund string model Andersson et al. (1983a), and the cluster fragmentation model Amati and Veneziano (1979). In this manuscript, we show that jet substructure selection can largely increase this sensitivity, allowing for more stringent tests to hadronization mechanisms. For that, this manuscript is organized as follows: in section II we will present an overview of the main hadronization models widely used by Monte Carlo hadronic event generators Pythia Bierlich et al. (2022) and Herwig-Bahr et al. (2008); the jet selection procedure to enhance hadronization effects, based on the resolved SoftDrop, will be presented in section III; results for the charge correlation ratio are discussed in section IV, with final conclusions in section V.
II II - Hadronization Models Overview
We generally expect that the evolution of a hard scattered parton ends with the parton attaining an on-shell mass for hadronization. A theoretical description of such process, i.e. the metamorphosis Ellis et al. (1976) of partons into hadrons, is unknown and is counted as one of the important unsolved problems in Physics. Hence, the treatment of hadronization has to reflect the non-perturbative conversion of the coloured, outgoing final-state partons into colourless hadrons through a combination of analytical results and different QCD-based phenomenological models. There are two main classes of hadronization models - the Lund string, implemented in Pythia, and the cluster fragmentation approaches, implemented in the Herwig and Sherpa Gleisberg et al. (2009) MC event generators.
The Lund String model Andersson et al. (1983a, b); Sjostrand (1984); Andersson (2023) is based on the assumption (supported by lattice QCD results Bali and Schilling (1992)) that the confining force field between a colour and an anticolour charge, as in a quark/antiquark colour-singlet $q\bar{q}$, can be accurately described as a flux tube with linearly-increasing potential energy with respect to the inter-charge distance $r$. This is accounted by the Cornell potential $V(r)=-\frac{4}{3}\frac{\alpha_{s}}{r}+k\,r$, originally proposed to describe quarkonium masses, where the Coulomb term becomes negligible for large $r$. In the linear confinement framework, as the extremities of a $q\bar{q}$ dipole move apart from their common production vertex, the stretching of the colour flux tube can be modeled by a massless relativistic string with string tension $k\approx 1$ GeV/fm and no transverse degrees of freedom. Depending on the energies of the original pair, it may evolve into a stable so-called “yoyo” mode for smaller production energies, identified as mesons with flavour inferred from their quark content, or into a sequence of string fragmentations, for higher energies. For the case of a high-energy $q\bar{q}$, the potential energy of the string increases with parton separation until it is high enough for a new parton/antiparton pair to be formed. This leads the string to be broken up by the formation of new $q\bar{q}$ colour-singlet systems. After successive fragmentation, the energy scale reduces until on-shell hadrons ($p_{\text{hadron}}^{2}=m_{\text{hadron}}^{2}$) are produced. In the Lund String model, this is the point at which the fragmentation process is assumed to end. This picture is expanded to include gluons in Pythia, allowing $q\bar{q}$ strings to stretch through a finite number of gluons.
The cluster hadronization framework is based on the colour “preconfinement” property Amati and Veneziano (1979). Partons perturbatively evolve from the hard scattering scale $Q$ down to $Q_{0}$, where quarks and gluons rearrange themselves into on-shell lumps of colour singlets - clusters - whilst not yet allowing on-shell hadrons to emerge. This intermediate stage of jet development is what is called pre-confinement. It is then assumed that confinement, describing the processes that bring the scale $Q_{0}$ down to hadron masses $O\left(m_{\pi},m_{p}\right)$, will convert these limited mass, colourless clusters into hadrons. This is where the cluster fragmentation model comes into play. Herwig’s implementation starts by splitting the gluons at the end of the parton shower into (di)quark/anti(di)quark pairs Bahr et al. (2008), constrained by the reminiscent decay phase-space. After this first step, the event is fully reduced to colour-connected quark/antiquark or diquark/antidiquark pairs, to be grouped in the model’s colour singlet clusters. The cluster mass spectrum is independent on the collision system and centre-of-mass energy and is peaked at low masses, meaning clusters can be taken as highly-excited hadron resonances and decayed, according to phase-space, into the observed hadrons. Nonetheless, the hadrons in the large-mass tail of the cluster spectrum require a cluster fission process to split them into lighter clusters before decaying. This is done by producing a $q\bar{q}$ pair from vacuum, forming two new clusters, each with one of the original partons. Once all clusters pass the model’s dynamical mass cutoff, the new set of clusters are finally decayed into pairs of on-shell hadrons.
III III - Event generation and jet substructure
@@
DIS events provide a clean environment to study the hadronization mechanism and confinement dynamics, as they produce final states with typically low multiplicities, allowing for more precise measurements. DIS data is also required for the testing and calibration of the theoretical tools, such as MC event generators, used to study the quark-gluon plasma produced in heavy-ion collisions. Here, we build on recent work Chien et al. (2022) related to electric charge and formation time studies in DIS systems towards refining the sensitivity of novel experimental observables to hadronization effects. As such, we make use of both Pythia 8.306 and Herwig 7 to test a variety of jet observables and selections across the two different phenomenological models - Lund and cluster, respectively - described in the previous Section. We aim to inform future experimental measurements at the upcoming Electron Ion Collider (EIC), planned to start its activity by the start of the new decade.
@@
As such, we simulate 40 million $ep$ DIS collisions using both MC event generators, anticipated to be 4% of the total integrated luminosity at EIC Chien et al. (2022). We impose the maximum EIC beam energies of $\sqrt{s_{e}}=18$ GeV and $\sqrt{s_{p}}=275$ GeV for the electron and proton beams, respectively. A minimum momentum transfer of $Q^{2}>50$ GeV2/c2 is applied on the hard scatterings to ensure the production of high $p_{T}$ jets and particles are selected with transverse momentum $p_{T}>200$ MeV/c to remove high-collinearity to the beam.
@@
From the list of final-state particles, jet finding via the anti-kT algorithm Cacciari et al. (2008) as implemented in fastjet 3.3.0 Cacciari et al. (2012) with a resolution parameter of $R=1$ is performed and inclusive $p_{T,jet}>7$ GeV/c [^1] While we closely followed the procedure depicted in Chien et al. (2022), we focused on $p_{T,jet}>7~{}\rm{GeV/c}$ instead of $p_{T,jet}>5~{}\rm{GeV/c}$ to minimize the differences between the hadronization models. transverse momentum jets are selected within a pseudo-rapidity range of $-3.5<\eta_{jet}<1.5$. We then perform the SD grooming on jets via the Cambridge/Aachen (C/A) algorithm Dokshitzer et al. (1997) for geometric substructure analysis, since it enforces a strict angular ordering in the clustering trees. Accordingly, the first declustering step of these trees corresponds to the largest opening angle. For formation time studies, the SD algorithm is applied to $\tau$ re-clustered trees Apolinário et al. (2021, 2024) since the $\tau$ algorithm’s distance measure becomes a proxy for the (inverse) formation time of the clustered particles in the high-energy, soft and collinear limits. As such, the first declustering will generally have the shortest formation time value.
@@
With the goal of comparing hadronization models in mind, we select the parton shower descriptions in Pythia and Herwig that produce the best matches between profiles of the event’s particles and jets, with respect to observables like azimuthal angle ($\phi$), pseudorapidity ($\eta$) and transverse momentum ($p_{T}$). For the chosen collision settings, the results matched the most for Pythia’s simple shower and Herwig’s dipole shower. For reference, Fig. 1 shows the particle (top) and jet (bottom) transverse momentum spectra for the DIS events, in full circles for Pythia simulations and in open circles for Herwig with the settings and selections mentioned above, along with the ratios with respect to Herwig results. We can see in the jet spectra that variations between the Monte Carlo generators fall below 10% for a wide range of transverse momenta.
@@
where the transverse momentum of two objects $(1,2)$ are numbered such that $p_{T,2}\leq p_{T,1}$. The SD criterion imposed on the C/A re-clustered trees follows the general selection
@@
Finally, we can also evaluate the formation time $\tau_{\text{form}}$ Apolinário et al. (2024), which is the time that a quantum state, such as a parton, takes to behave as two independent sources of additional radiation, for example, the two daughter-partons from a splitting. The formation time within jets is associated to varying stages of the parton evolution, ranging from pQCD to npQCD effects. For a jet splitting, $\tau_{\text{form}}$ can be calculated by:
@@
Since vacuum parton showers typically generate angular ordered emissions and the re-clustering procedure is performed with an angular ordering variable via the Cambridge/Aachen algorithm, the first unclustering step to survive the SD criterion will produce prongs with larger $R_{g}$ within the jet (proxy for the opening angle), followed by a sequential reduction in the angle as we travel along the clustering tree. This is reflected on the top panel from Fig. 3, where the 1SD, in the blue markers (full circles for Pythia simulations and open circles for Herwig), will be on average observed at larger angles ($R_{g,1SD}\approx 0.52/0.56$ for Pythia/Herwig). Since the RSD, by construction, happens either at or after the 1SD in the clustering tree, their opening angles in the black markers are shown to be smaller ($R_{g,RSD}\approx 0.33/0.40$). As for the LCP, represented via the blue markers, we find that it closely follows the $R_{g}$ distribution of the RSD splitting, meaning the subsequent branching of the RSD daughters will mostly preserve the opening angle between the leading charged particles ($R_{g,LCP}\approx 0.33/0.40$). The $R_{g}$ is shown to be sensitive to the hadronization procedure, with significant differences between Pythia and Herwig.
@@
IV IV - Formation Time and Charge Correlation Ratio
@@
With RSD focusing on the high-energy charged particles within the jet, we can directly compare this study with the charge-correlation ratio observable introduced here Chien et al. (2022). This is defined as the ratio of yields of equally-charged, $d\sigma_{hh}/dX$, and oppositely-charged, $d\sigma_{h\overline{h}}/dX$, leading charged hadrons of flavour $h=\pi^{+},K,p$ (pions, kaons and protons), within the jets
@@
Therefore, $r_{c}$ acts as a measure of the probability of producing jets with a given LCP charge configuration, with values ranging between $11\leq r_{c}\leq-$ 1: positive $r_{c}$ means a greater probability of jets having both positive or both negative LCP (same sign), while negative $r_{c}$ signals that jets tend to have both a positive and a negative leading charges (opposite sign). This was recently shown as a direct test of the string breaking phenomenon.
Since the definition of the $r_{c}$ is meant to be sensitive to hadronization via the two hardest hadrons, we can study it as a function of the formation time defined in Eq. (4). The charge ratio dependence of the LCP formation time for jets with leading pions, kaons and protons is shown in Fig. 5  Chien et al. (2022), in the black, red and blue markers, respectively, sampled from jets with at least two identified charged particles. The analysis of Pythia(Herwig) jets, corresponding to the charge ratios in full(open) circle markers, reveals that 55%(57%) have a leading pair of charged pions, 5%(3%) have leading kaons and 4%(6%) have leading protons.
@@
Firstly, the charge ratios are shown to be overall negative, consistently across hadron species. However, the $r_{c}$ is clearly dependent on the formation time of the LCP. For large $\tau_{\text{form, LCP}}$, the charge ratios are approximately constant at large negative values of $r_{c}$, particularly for $KK$ and $pp$ pairs within jets. In fact, jets with leading protons have a near 100% probability of having $p\bar{p}$, with highly unlikely $pp$ and $\bar{p}\bar{p}$ configurations. This reveals a strong preference towards the production of opposite leading charged particles for late LCP jets. We recall that RSD is a good proxy for the LCP for large $\tau_{\text{form, LCP}}$ and that $\tau$ re-clustered trees will have the largest formation time values for the latest unclustering steps, i.e., in the last stages of the jet fragmentation pattern. As such, large LCP formation time selects jets whose leading charged hadrons get resolved very late, making them less likely to have subsequent branching randomizing the correlation between the charges.
@@
IV.1 Resolved SoftDrop depth
Our goal is to increase the sensitivity of this charge ratio variable to hadronization physics. Namely, jets can have their resolved SD earlier or later in their respective fragmentation pattern. When selecting jets with an identifiable RSD splitting that resolves the leading charged particles from that jet, we observe $\langle N_{SD}\rangle\approx 3$ SoftDrop unclusterings along the main branch for both for Pythia and Herwig, as can be seen in Fig. 6 on the top and bottom panels, respectively. As such, the contributions coming from jets with RSD in the first, second and third SDs are shown as stacked $N_{SD}$ distributions in red, blue and green bars, respectively. We find that 52% of Pythia jets and 59% of Herwig jets have their resolved splitting at the very first SD unclustering of the clustering tree. The RSD will fall on the second SD for an additional 30%/26% of jets and on the third SD for 13%/11% of them for the kinematic range studied.

### v3

# Enhancing charge ratio sensitivity to hadronization effects via jet selections on resolved SoftDrop splitting

@@ Abstract
The study of Quantum Chromodynamics (QCD) at ultra-relativistic energies can be performed in a controlled environment through lepton-hadron deep inelastic scatterings. In such collisions, the high-energy partonic emissions that follow from the ejected hard partons are accurately described by perturbative QCD. However, the lower energy scales at which quarks and gluons experience colour confinement, i.e. hadronization mechanism, fall outside the validity regions for perturbative calculations, requiring phenomenological models tuned to data to describe it. As such, hadronization physics cannot be currently derived from first principles alone. Monte Carlo event generators are useful tools to describe these processes as they simulate both the perturbative and the non-perturbative interactions, with model-dependent energy scales that control parton dynamics. This work employs jets - experimental reconstructions of final-state particles likely to have a common partonic origin - to inspect this transition further. Although originally proposed to circumvent hadronization effects, we show that jets can be utilised as probes of non-perturbative phenomena via their substructure. The charge correlation ratio was recently shown to be sensitive to hadronization effects. Our work further improves this sensitivity to non-perturbative scales by introducing a new selection based on the relative placement of the resolved SoftDrop splitting within the clustering tree, defined as the unclustering that resolves the jet’s leading charged particles.
@@ I I - Introduction
Their treatment of hadronization has to reflect the non-perturbative conversion of the coloured, outgoing final-state partons into colourless hadrons through a combination of analytical results and different QCD-based phenomenological models. There are two main classes of hadronization models - the Lund string model Andersson et al. (1983a, b); Sjostrand (1984); Andersson (2023); Bali and Schilling (1992), implemented for instance in Pythia Bierlich et al. (2022), and the cluster fragmentation approach Amati and Veneziano (1979), found in Herwig Bahr et al. (2008) and Sherpa Gleisberg et al. (2009) MC event generators. While the Lund string model depicts hadronization after the fragmentation of massless, one-dimensional strings stretching between colour-connected charges, the cluster hadronization framework is based on the colour “preconfinement” property and prescribes the fragmentation of on-shell, colour singlet lumps of quarks and gluons - the clusters. For more information, we refer the reader for the references above.
@@
Recent effort in the high-energy physics community has been in the area of developing novel experimental algorithms that translate a jet clustering tree to a theoretically motivated description of a parton shower Dasgupta and Salam (2001); Dasgupta et al. (2013); Larkoski et al. (2014); Mehtar-Tani et al. (2020). Measurements from the Relativistic Heavy-Ion Collider (RHIC) and from the Large Hadron Collider (LHC) have shown that jet substructure observables, when calculated via grooming techniques such as SoftDrop (SD) Sirunyan et al. (2018a, b); Aad et al. (2020a); Acharya et al. (2020); Aad et al. (2020b); Adam et al. (2020); Abdallah et al. (2021); Acharya et al. (2022); ALICE (2022a, b), allow one to fit pQCD calculations at the first hard splitting without the need for large hadronization corrections. Namely, the groomed momentum fraction ($z_{g}$), i.e. the fractional transverse momentum of the softest subjet in a QCD jet, fits very well to the pQCD splitting functions when evaluated at the first SD splitting. For subsequent emissions along the higher energy (hard) branch of a jet shower Dreyer et al. (2018), the shape of the splitting function gradually changes from a steeply falling pQCD-like $1/z_{g}$ distribution to a flatter shape due to phase space restrictions. However, so as the impact of npQCD effects Robotkova (2022); Kunnawalkam Elayavalli (2022). Previous works Apolinário et al. (2021) demonstrated that SD groomed clustering trees significantly improve the correlation between the formation time of the parton shower emissions and the formation time of the first jet SD unclustering step, regardless of choice of jet clustering algorithm Apolinário et al. (2021). However, subsequent unclustering steps also observe their parton shower correlation worsen, although with a large dependence on the chosen algorithm.
@@
The current state of SoftDrop substructure studies therefore requires the assessment of non-perturbative contributions to the jet evolution history, as they become more and more meaningful to the late-time jet dynamics. This calls for the exploration of experimentally robust jet substructure observables sensitive to hadronization physics. In Chien et al. (2022) the charge correlation ratio was proposed as a possible observable that is able to discriminate between the two aforementioned hadronization models: Lund string and cluster fragmentation. In this manuscript, we show that jet substructure selection can largely increase this sensitivity, allowing for more stringent tests to hadronization mechanisms. For that, this manuscript is organized as follows: in section II we present the details of event generation using Pythia and Herwig and the jet selection procedure to enhance hadronization effects, based on the resolved SoftDrop; results for the charge correlation ratio are discussed in section III, with final conclusions in section IV. Finally, Appendix A discusses more in detail the impact of selecting different parton showers and hadronization models on the present study.
II II - Event generation and jet substructure
@@
DIS events provide a clean environment to study the hadronization mechanism and confinement dynamics, as they produce final states with typically low multiplicities, allowing for more precise measurements. DIS data is also required for the testing and calibration of the theoretical tools, such as MC event generators, used to study the quark-gluon plasma produced in heavy-ion collisions. Here, we build on recent work Chien et al. (2022) related to electric charge and formation time studies in DIS systems towards refining the sensitivity of novel experimental observables to hadronization effects. As such, we make use of both Pythia 8.306 and Herwig 7.2 to test a variety of jet observables and selections across the two different phenomenological models - Lund and cluster, respectively. We aim to inform future experimental measurements at the upcoming Electron Ion Collider (EIC), planned to start its activity by the start of the new decade.
@@
As such, we simulate 40 million $ep$ DIS collisions using both MC event generators, anticipated to be 4% of the total integrated luminosity at EIC Chien et al. (2022). We impose the maximum EIC beam energies of $E_{e}=18$ GeV and $E_{p}=275$ GeV for the electron and proton beams, respectively. A minimum momentum transfer of $Q^{2}>50$ GeV2/c2 is applied on the hard scatterings to ensure the production of high $p_{T}$ jets and particles are selected with transverse momentum $p_{T}>200$ MeV/c to remove high-collinearity to the beam.
@@
From the list of final-state particles, jet finding via the anti-kT algorithm Cacciari et al. (2008) as implemented in fastjet 3.3.0 Cacciari et al. (2012) with a resolution parameter of $R=1$ is performed and inclusive $p_{T,jet}>7$ GeV/c [^1] While we closely followed the procedure depicted in Chien et al. (2022), we focused on $p_{T,jet}>7~{}\rm{GeV/c}$ instead of $p_{T,jet}>5~{}\rm{GeV/c}$ to minimize the differences between the hadronization models. transverse momentum jets are selected within a pseudo-rapidity range of $-3.5<\eta_{jet}<1.5$. As we intend to probe hadronization using jets, specifically by studying the electric charges of their highest transverse momentum particles, only jets with at least two final-state charged hadrons are selected. We then perform the SD grooming on jets via the Cambridge/Aachen (C/A) algorithm Dokshitzer et al. (1997) for geometric substructure analysis, since it enforces a strict angular ordering in the clustering trees. Accordingly, the first unclustering step of these trees corresponds to the largest opening angle. For formation time studies, the SD algorithm is applied to $\tau$ reclustered trees Apolinário et al. (2021, 2024) since the $\tau$ algorithm’s distance measure becomes a proxy for the (inverse) formation time of the clustered particles in the high-energy, soft and collinear limits. As such, the first unclustering will generally have the shortest formation time value.
@@
With the goal of comparing hadronization models in mind, we select the parton shower descriptions in Pythia and Herwig that produce the best matches between profiles of the event’s particles and jets, with respect to observables like azimuthal angle ($\phi$), pseudorapidity ($\eta$) and transverse momentum ($p_{T}$). For the chosen collision settings, the results matched the most for Pythia’s simple shower and Herwig’s dipole shower. For reference, Fig. 1 shows the particle (top) and jet (bottom) transverse momentum spectra for the DIS events, in full circles for Pythia simulations and in open circles for Herwig with the settings and selections mentioned above, along with the ratios with respect to Herwig results. We can see in the jet spectra that variations between the Monte Carlo generators fall below 10% for a wide range of transverse momenta. [^2] Check Appendix A for more details on the effects of the parton shower and hadronization model.
@@
where the transverse momentum of two objects $(1,2)$ are numbered such that $p_{T,2}\leq p_{T,1}$. The SD criterion imposed on the C/A or $\tau$ reclustered trees follows the general selection
@@
Finally, we can also evaluate the formation time $\tau_{\text{form}}$ Apolinário et al. (2021, 2024), which is the time that a quantum state, such as a parton, takes to behave as two independent sources of additional radiation, for example, the two daughter-partons from a splitting. The formation time within jets is associated to varying stages of the parton evolution, ranging from pQCD to npQCD effects. For a jet splitting, $\tau_{\text{form}}$ can be calculated by:
@@
Since vacuum parton showers typically generate angular ordered emissions and the re-clustering procedure is performed with an angular ordering variable via the Cambridge/Aachen algorithm, the first unclustering step to survive the SD criterion will produce prongs with larger $R_{g}$ within the jet (proxy for the opening angle), followed by a sequential reduction in the angle as we travel along the clustering tree. This is reflected on the top panel from Fig. 3, where the 1SD, in the blue markers (full circles for Pythia simulations and open circles for Herwig), will be on average observed at larger angles ($R_{g,1SD}\approx 0.52/0.56$ for Pythia/Herwig). Since the RSD, by construction, happens either at or after the 1SD in the clustering tree, their opening angles in the black markers are shown to be smaller ($R_{g,RSD}\approx 0.32/0.36$). As for the LCP, represented via the blue markers, we find that it closely follows the $R_{g}$ distribution of the RSD splitting, meaning the subsequent branching of the RSD daughters will mostly preserve the opening angle between the leading charged particles ($R_{g,LCP}\approx 0.32/0.36$). The $R_{g}$ is shown to be sensitive to the hadronization procedure, with significant differences between Pythia and Herwig.
@@
III III - Formation Time and Charge Correlation Ratio
@@
With RSD focusing on the high-energy charged particles within the jet, we can directly compare this study with the charge correlation ratio observable introduced here Chien et al. (2022). This is defined as the ratio of yields of equally-charged, $d\sigma_{hh}/dX$, and oppositely-charged, $d\sigma_{h\overline{h}}/dX$, leading charged hadrons of flavour $h=\pi^{+},K,p$ (pions, kaons and protons), within the jets
@@
Therefore, $r_{c}$ acts as a measure of the probability of producing jets with a given LCP charge configuration, with values ranging between $-1\leq r_{c}\leq 1$: positive $r_{c}$ means a greater probability of jets having both positive or both negative LCP (same sign), while negative $r_{c}$ signals that jets tend to have both a positive and a negative leading charges (opposite sign). This was recently shown as a direct test of the string breaking phenomenon.
Since the definition of the $r_{c}$ is meant to be sensitive to hadronization via the two hardest hadrons, we can study it as a function of the formation time defined in Eq. (4). The charge ratio dependence on the LCP formation time for jets with leading pions, kaons and protons is shown in Fig. 5  Chien et al. (2022), in the black, red and blue markers, respectively, sampled from jets with at least two identified charged particles. The analysis of Pythia(Herwig) jets, corresponding to the charge ratios in full(open) circle markers, reveals that 61%(66%) have a leading pair of charged pions, 5%(3%) have leading kaons and 5%(6%) have leading protons, with the remaining jets having mixed-flavour LCP.
@@
Firstly, the charge ratios are shown to be overall negative, consistently across hadron species. However, the $r_{c}$ is clearly dependent on the formation time of the LCP. For large $\tau_{\text{form, LCP}}$, the charge ratios are approximately constant at large negative values of $r_{c}$, particularly for $KK$ and $pp$ pairs within jets. In fact, jets with leading protons have a near 100% probability of having $p\bar{p}$, with highly unlikely $pp$ and $\bar{p}\bar{p}$ configurations. This reveals a strong preference towards the production of opposite leading charged particles for late LCP jets. We recall that RSD is a good proxy for the LCP for large $\tau_{\text{form, LCP}}$ and that $\tau$ reclustered trees will have the largest formation time values for the latest unclustering steps, i.e., in the last stages of the jet fragmentation pattern. As such, large LCP formation time selects jets whose leading charged hadrons get resolved very late, making them less likely to have subsequent branching randomizing the correlation between the charges.
@@
III.1 Resolved SoftDrop depth
Our goal is to increase the sensitivity of this charge ratio variable to hadronization physics. Namely, jets can have their resolved SD earlier or later in their respective fragmentation pattern. When selecting jets with an identifiable RSD splitting that resolves the leading charged particles from that jet, we observe $\langle N_{SD}\rangle\approx 3$ SoftDrop unclusterings along the main branch for both for Pythia and Herwig, as can be seen in Fig. 6 on the top and bottom panels, respectively. As such, the contributions coming from jets with RSD in the first, second and third SDs are shown as stacked $N_{SD}$ distributions in red, blue and green bars, respectively. We find that 52% of Pythia jets and 55% of Herwig jets have their resolved splitting at the very first SD unclustering of the clustering tree. The RSD will fall on the second SD for an additional 30%/28% of jets and on the third SD for 13%/12% of them for the kinematic range studied.